OpenStudy (anonymous):

determine the slopeof the line giveneach pair of points (-1 , -1 ) and (5 , -3)

OpenStudy (anonymous):

how to solve it?

OpenStudy (anonymous):

typo! \[\Large \implies m=\frac{-3-(-1)}{5-(-1)}=\frac{-2}{6}=\frac{-1}{3}\]

OpenStudy (anonymous):

@alloids65432 got it?

OpenStudy (anonymous):

oh ty

OpenStudy (anonymous):

yw:)

OpenStudy (anonymous):

yw?

OpenStudy (anonymous):

you 're welcome:)

OpenStudy (anonymous):

ahh kk thnx

OpenStudy (anonymous):

1 more question

OpenStudy (anonymous):

how about if it is like this (-2,5) and (-2,8) ??

OpenStudy (anonymous):

ok just remember the following slope=difference of y values/difference of x values \[\Large m=\frac{8-5}{-2-(-2)}\]